the bases of a trapezoidal field are 225 meters and 125 meters, if the distance between them is ninety-two meters, find its area in hectares
Answer by fcabanski(1391)

The area of a trapezoid is 1/2 * (b1+b2)*h where b1 and b2 are the lengths of the bases and h is the height or distance between the bases.


A = 1/2 * (225+125)*92 = 16100 square meters.


There are 10,000 square meters in a hectare.


16100/10000 = 1.61 hectares.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com